OpenStudy (anonymous):

I Need Help With This Question What Is The Graph Of The Function Rule? y = x + 1

OpenStudy (tkhunny):

A line?

OpenStudy (anonymous):

@tkhunny I have no clue

OpenStudy (anonymous):

that's the question I don't no how to do the problem

OpenStudy (tkhunny):

We are going to need a better question. There don't appear to be any instructions. Shall we graph it? Talk about it? Buy it roses and take it to the prom? What?

OpenStudy (anonymous):

lol!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! #DEAD

OpenStudy (anonymous):

BUT I THINK IM POST TO GRAPH IT THO

OpenStudy (tkhunny):

We can graph it. Have you met the term "Slope-Intercept Form"? You may have seen an example like this, "y = mx + b".

OpenStudy (anonymous):

YES

OpenStudy (tkhunny):

Excellent. We have this: y = x+1 This is already in slope-intercept form. m = 1 (the invisible coefficient on the x) b = 1 (the constant term) Agreed?

OpenStudy (anonymous):

AGREED

OpenStudy (tkhunny):

The y-intercept is 1, making a point on the graph, (0,1). Plot this point and we are well on our way.

OpenStudy (anonymous):

So I just the graph and plot (0,1)?

OpenStudy (tkhunny):

That's where we start on this one. The slope-intercept form gives us that y-intercept. We stll need more information. That's just the place to start.

OpenStudy (anonymous):

ok ive graphed it

OpenStudy (tkhunny):

Now we need the slope. In this case, m = 1. A slope of 1. "Slope" is an odd sort of thing. You do have to get used to it a little. You may have heard "rise over run" or "change in y divided by change in x". These are good descriptions, but a little abstract. I prefer to think of it in two pieces... 1) You PROMISE to move 1 space in the positive x direction. 2) How far in the positive y direction shall I move? It is the agreement in #2 that is the slope. x always moves 1 and y always moves "the slope". With a slope of 1, this means as always, x moves one to the right and y moves 1 (the slope) up. This should bring us to the point (1,2) -- one right and slope up.

OpenStudy (anonymous):

so I graph (1,2)?

OpenStudy (tkhunny):

That's where the slope took us from (0,1). +1 for x and +Slope for y. Then we're done. Two points determine a line, so get a really long ruler, connect the dots, and extend the line as far as you can in both directions.

OpenStudy (anonymous):

so I was doing something but im back

OpenStudy (anonymous):

ok so let me get this straight I stayed on the line because of the "0" and I went up "1" and put my dot then now I stay on the line and I go over "1" and up "2"? @tkhunny

OpenStudy (tkhunny):

y-intercept of 1 says start at (0,1) Slope of 1 says move to (0+1,1+Slope) = (1,2)

OpenStudy (anonymous):

so Im right? there should be two dots connecting and a long line?

OpenStudy (tkhunny):

I'm nervous about this: " go over "1" and up "2"" You should go over 1 and up 1 (the slope) and this should get you to (1,2). So, you MAY be saying the right thing. I just can't tell.

OpenStudy (anonymous):

but I graphed the the (0,1) already?????????????

OpenStudy (tkhunny):

What's fine. What other point did you graph, based on what the slope told you to do?

OpenStudy (anonymous):

(0,1) & (1,2) ???????? and got two dots and straight line?

OpenStudy (tkhunny):

Those are the right two points. Please draw a line through them and be done!

OpenStudy (anonymous):

ok perfect but one quick question where and why did I have to go over 2?

OpenStudy (anonymous):

@tkhunny ^^^

OpenStudy (tkhunny):

I never said that. That's why I made you repeat it when you said it. It was incorrect. The slope is 1. This means x moves 1 and y moves 1. No 2.

OpenStudy (anonymous):

so what the correct graphing?

OpenStudy (tkhunny):

You MUST pay better attention. This is the third time we have concluded this problem. The y-intercept say plot (0,1). The slope says plot (1,2) This defines the line to graph. Connect the two points and you are done.

OpenStudy (anonymous):

are u changing the x into the number one and adding it? because the problem only gave the number 1 and I have already graphed it that way I just want to make sure this sticks in my head @tkhunny

OpenStudy (tkhunny):

I do not know what else to tell you. There are only three pieces. They are clearly defined in my last post.

OpenStudy (anonymous):

got ya thanks for the help

OpenStudy (anonymous):

@tkhunny I got the problem down but I really need help with this next one

OpenStudy (tkhunny):

If it's in slope-intercept form, yo uneed to show me ALL of the contrsuction. Name the three pieces.

OpenStudy (anonymous):

it isn't in slop-intercept form @tkhunny

OpenStudy (anonymous):

I bumped it up to the top @tkhunny